two cars are moving with same velocity c/2 .these are perpendicular to each other .find relative velocity.
1
Expert's answer
2020-09-14T10:23:54-0400

As Lorenz transformation is "u=\frac{\sqrt{(u'^2+v^2+2vu'\cos\theta'-(\frac{vu'\sin\theta'}{c})^2}}{1+\frac{vu'\cos\theta'}{c^2}}" (J.D.Jackson,Classical Electrodynamics, N.Y-London,1962)

We have "u'=v=c/2, \theta'=\pi/2" then "u=c\sqrt{7/16}"